Grade 11Mechanics

A boat is sent across a river with a velocity of 8 km/h .If the resultant velocity of boat is 10 km/he then velocity of river water is

Priyanka Dabade
9 Years agoGrade 11

2 Answers

Vidura
9 Years ago
The boat’s velocity and the river velocity will give the boat a new resultant velocity.
As the boat’s velocity is perpendicular to river velocity,
 
Assume that,
                   V= Velocity of boat
                   VR = Velocity of river
 
Apply pythagoras theoram,
 
Resultant velocity2 =  VB2 + VR2
                       102 =  82 + VR2
                               100 = 64 + VR2
                               VR2  = 36
                       VR  = 6 km/h
Rishi Sharma
5 Years ago
Dear Student,
Please find below the solution to your problem.

The boat’s velocity and the river velocity will give the boat a new resultant velocity.
As the boat’s velocity is perpendicular to river velocity,

Assume that,
VB= Velocity of boat
VR = Velocity of river

Apply pythagoras theoram,

Resultant velocity^2 = VB^2+VR^2
10^2=8^2+VR^2
100 = 64 +VR^2
VR^2 = 36
VR = 6 km/h

Thanks and Regards